OpenStudy (anonymous):

Hannah earned $1536 in 8 years on an investment at a 4% simple interest rate. How much was Hannah's investment? Hannah earned $1536 in 8 years on an investment at a 4% simple interest rate. How much was Hannah's investment? @Mathematics

OpenStudy (anonymous):

4% a year with simple interest for 8 years = 8 * 4% = 32%. 32% of Investment = 1536. Investment = 1536/32% = 1536 * 100 / 32 = ?

OpenStudy (anonymous):

4800 if your too lazy to add